### v2

# LLM4Mat-Bench: Benchmarking Large Language Models for Materials Property Prediction

Abstract
Large language models (LLMs) are increasingly being used in materials science. However, little attention has been given to benchmarking and standardized evaluation for LLM-based materials property prediction, which hinders progress. We present LLM4Mat-Bench, the largest benchmark to date for evaluating the performance of LLMs in predicting the properties of crystalline materials. LLM4Mat-Bench contains about 1.9M crystal structures in total, collected from 10 publicly available materials data sources, and 45 distinct properties. LLM4Mat-Bench features different input modalities: crystal composition, CIF, and crystal text description, with 4.7M, 615.5M, and 3.1B tokens in total for each modality, respectively. We use LLM4Mat-Bench to fine-tune models with different sizes, including LLM-Prop and MatBERT, and provide zero-shot and few-shot prompts to evaluate the property prediction capabilities of LLM-chat-like models, including Llama, Gemma, and Mistral. The results highlight the challenges of general-purpose LLMs in materials science and the need for task-specific predictive models and task-specific instruction-tuned LLMs in materials property prediction  [^1] The Benchmark and code can be found at: https://github.com/vertaix/LLM4Mat-Bench.
Keywords: large language models, materials property prediction, crystalline materials, benchmarks
1 Introduction
With the remarkable success of large language models (LLMs) in solving natural language tasks (Devlin et al., 2018; Radford et al., 2018; Raffel et al., 2020; Achiam et al., 2023; Touvron et al., 2023) and different scientific tasks (Lin et al., 2022; Edwards et al., 2022; Valentini et al., 2023; Castro Nascimento and Pimentel, 2023; Fang et al., 2023; Lv et al., 2024), scientists have recently started to leverage LLMs to tackle very important and challenging problems in materials science, including predicting materials properties (Rubungo et al., 2023; Korolev and Protsenko, 2023; Xie et al., 2023; Das et al., 2023; Qu et al., 2024; Choudhary, 2024) and discovering new materials (Antunes et al., 2023; Gruver et al., 2023; Qu et al., 2024; Choudhary, 2024).
The learning capabilities of LLMs have the potential to revolutionize the field of materials science. For example, recent research by Rubungo et al. (2023) has demonstrated the exceptional performance of LLMs in predicting the properties of crystalline materials based on textual descriptions of their structures. In their study, they introduced a novel dataset, TextEdge, which comprises textual descriptions of crystals and their corresponding properties. This dataset was used to fine-tune the encoder component of the T5-small model for the task of materials property prediction. The findings of Rubungo et al. (2023) challenge the conventional practice of heavily relying on graph neural networks and using solely either crystal composition or structure as input for property prediction. Their work underscores the significance of further investigating the extent to which LLMs can be harnessed to develop innovative techniques for accurately predicting the properties of crystalline materials, thereby enhancing the materials discovery pipeline. Unfortunately, the proposed TextEdge dataset is limited in scope, comprising approximately 145K samples and encompassing only three distinct properties. Furthermore, its lack of diversity, being derived from a single data source (Materials Project (Jain et al., 2013)), hinders its effectiveness in assessing the robustness of LLMs in materials property prediction.
In this work, we introduce LLM4Mat-Bench, a benchmark dataset collected to evaluate the performance of LLMs in predicting the properties of crystalline materials. To the best of our knowledge, LLM4Mat-Bench is the most extensive benchmark to date for assessing the efficacy of language models in materials property prediction. The dataset comprises approximately two million samples, sourced from ten publicly available materials sources, each containing between 10K and 1M structure samples. LLM4Mat-Bench encompasses several tasks, including the prediction of electronic, elastic, and thermodynamic properties based on a material’s composition, crystal information file (CIF), or textual description of its structure. We use LLM4Mat-Bench to evaluate several LLMs of different sizes, namely LLM-Prop (Rubungo et al., 2023) (35M parameters), MatBERT (Walker et al., 2021) (109.5M parameters), and Llama2 (Touvron et al., 2023) (7B parameters). And we provide fixed train-valid-test splits, along with carefully designed zero-shot and few-shot prompts to ensure reproducibility. We anticipate that LLM4Mat-Bench will significantly advance the application of LLMs in addressing critical challenges in materials science, including property prediction and materials discovery.
2 LLM4Mat-Bench
2.1 Data Collection Process
We collected the data used to create LLM4Mat-Bench from 10 publicly available materials data sources. In this section, we describe each data source and discuss how we accessed its data.
2.1.1 Data sources
hMOF (Wilmer et al., 2012) is a publicly available database [^2] https://mof.tech.northwestern.edu/ consisting of about 160K Metal-Organic Frameworks (MOFs), generated by Wilmer et al. using computational approaches. Materials Project (MP) (Jain et al., 2013) is a database with around 150K materials, offering free API access [^3] https://next-gen.materialsproject.org/api for data retrieval, including CIF files and material properties. The Open Quantum Materials Database (OQMD) (Kirklin et al., 2015) is a publicly accessible database [^4] https://www.oqmd.org/ of 1.2M materials, containing DFT-calculated thermodynamic and structural properties, created at Northwestern University. OMDB (Borysov et al., 2017) is an organic materials database with about 12K structures and related electronic band structure properties, freely available [^5] https://omdb.mathub.io/. JARVIS-DFT (Choudhary et al., 2017; Choudhary et al., 2018) is a repository created by NIST researchers, containing around 75.9K material structures with downloadable properties [^6] https://jarvis.nist.gov/jarvisdft. QMOF (Rosen et al., 2021, 2022) is a quantum-chemical property database of over 16K MOFs, accessible via GitHub [^7] https://github.com/Andrew-S-Rosen/QMOF. JARVIS-QETB (Garrity and Choudhary, 2023) is a NIST-created database [^8] https://jarvis.nist.gov/jarvisqetb/ of nearly one million materials with tight-binding parameters for 65 elements. GNoME is a database of 381K new stable materials discovered by Merchant et al. (2023) using graph networks and DFT, available on GitHub [^9] https://github.com/google-deepmind/materials_discovery/blob/main/DATASET.md. Cantor HEA (Li et al., 2024) is a DFT dataset of formation energies for 84K alloy structures, available on Zenodo [^10] https://doi.org/10.5281/zenodo.10854500. SNUMAT is a database with around 10K experimentally synthesized materials and DFT properties, accessible via API [^11] https://www.snumat.com/apis.
2.1.2 Collecting crystal information files (CIFs) and materials property
Crystal structure files (CIFs), material compositions, and material properties were collected from publicly accessible sources described in Section 2.1.1. Data collection was facilitated by APIs and direct download links provided by the respective databases. For databases such as Materials Project, OMDB, SNUMAT, JARVIS-DFT, and JARVIS-QETB, user registration is required for access, while databases like hMOF, QMOF, OQMD, and GNoME allow direct data access without registration. From each source, we obtained CIFs and associated material properties. Although the Materials Project and JARVIS-DFT databases offer a broader range of properties, we selected a subset—10 and 20 properties respectively—that adequately represents the data within our benchmark, based on the number of data points available for each property. This selection was made to optimize computational efficiency when training models across the 65 properties included in LLM4Mat-Bench.
2.1.3 Generating the textual description of crystal structure
LLMs perform better with textual input, and Rubungo et al. (2023); Korolev and Protsenko (2023); Qu et al. (2024) have demonstrated that LLMs can effectively learn the structural representation of a crystal from its textual description, outperforming graph neural network (GNN)-based models that directly utilize the crystal structure for property prediction.
Crystal structures are typically described in file formats such as Crystallographic Information File (CIF) which include predominantly numbers describing lattice vectors and atomic coordinates and are less amenable to LLMs. Instead of directly using these as inputs, we use Robocrystallographer (Jain et al., 2013) to deterministically generate texts that are more descriptive of crystal structures from CIF files. Robocrystallographer was developed and has been used by the Materials Project team to auto-generate texts for their database. Given a structure, Robocrystallographer leverages predefined rules and existing libraries to extract chemical and structural information, including oxidation states, global structural descriptions (symmetry information, prototype matching, structural fingerprint calculations etc.), and local structural descriptions (e.g. bonding and neighbor analysis, connectivity). This method not only generate deterministic and human-readable texts, but also ensures no data contamination in our fine-tuned LLMs, as the data sources mentioned do not include these crystal text descriptions.
Table 1: LLM4Mat-Bench statistics.
Data source # Structure files # Structure-Description pairs # Tokens (Words) # Avg. subword tokens/Sample # Properties
Total Train Validation Test Composition Structure Description Composition Structure Description
OQMD (Kirklin et al., 2015) 1,008,266 964,403 771,522 96,440 96,441 964K 96M 244M 5.3 635.4 347.3 2
JARVIS-QETB (Garrity and Choudhary, 2023) 829,576 623,989 499,191 62,399 62,399 624K 45M 90M 3.5 466.6 202.5 4
GNoME (Merchant et al., 2023) 381,000 376,276 301,020 37,628 37,628 830K 78M 508M 9.7 1185.3 1711.3 6
Materials Project (Jain et al., 2013) 146,143 146,143 125,825 10,000 10,318 272K 37M 157M 6.8 1611.8 1467.3 10
hMOF (Wilmer et al., 2012) 133,524 132,743 106,194 13,274 13,275 449K 96M 581M 14.9 4583.9 5629.3 7
Cantor HEA (Li et al., 2024) 84,024 84,019 67,215 8,402 8,402 84K 11M 251M 9.5 868.4 4988.6 4
JARVIS-DFT (Choudhary et al., 2017; Choudhary et al., 2018) 75,965 75,965 60,772 7,596 7,597 76K 9M 25M 5.0 786.0 455.9 20
QMOF (Rosen et al., 2021, 2022) 16,340 7,656 6,124 766 766 8K 7M 22M 14.0 5876.4 3668.0 4
OMDB (Borysov et al., 2017) 12,500 12,122 9,697 1,212 1,213 66K 8M 14M 14.8 4097.4 1496.6 1
SNUMAT  [^12] https://www.snumat.com/ 10,481 10,372 8,297 1,037 1,038 16K 2M 4M 5.9 1244.5 539.1 7
Total 2,697,779 1,978,985 1,592,315 193,357 193,313 4.7M 615.5M 3.1B 7.9 1559.7 1703.6 65
Table 2: Comparing the LLM4Mat-Bench with other existing benchmarks.
Benchmark # Data Sources # Distinct Properties # Properties/# Samples # Properties/Task Type Material Representations
<10k 10-100k 100k+ Regression Classification Composition Structure Description
MatBench (Dunn et al., 2020) 6 10 7 3 3 10 3
TextEdge (Rubungo et al., 2023) 1 3 0 0 3 2 1
LLM4Mat-Bench (Ours) 10 45 5 31 29 60 5
2.2 Data Statistics
As Table 1 shows, LLM4Mat-Bench comprises 2,697,779 structure files, which, after pairing with descriptions generated by Robocrystallographer and filtering out descriptions with fewer than five words, result in 1,978,985 composition-structure-description pairs [^13] The total number of pairs were 2,433,688, after removing about 454703 duplicated pairs across datasets, it resulted to 1,978,985 pairs.. The reduction in sample count is also due to Robocrystallographer’s inability to describe certain CIF files. The total samples for each dataset in LLM4Mat-Bench are randomly split into 80%, 10%, and 10% for training, validation, and testing, respectively. OQMD has the highest number of samples at 964,403, while QMOF has the fewest with 7,656 samples. On average, each dataset in LLM4Mat-Bench contains approximately 200,000 samples.
In LLM4Mat-Bench, when combined, textual descriptions contain 3.1 billion tokens, crystal structures 615 million, and compositions 4.7 million [^14] We used NLTK toolkit as a tokenizer to count the number of words/tokens.. OQMD leads in composition tokens (964K), while hMOF has the most description tokens (581M). For CIFs, both OQMD and hMOF have around 96M tokens. On average, compositions have 8 subword tokens per sample, CIFs 1600, and descriptions 1700. hMOF averages the longest inputs for compositions (14.9) and descriptions (5629), while QMOF leads in structures (5876.4) [^15] We used Llama 2 tokenizer to count the number of subword tokens.. JARVIS-DFT has the most tasks with 20 properties, followed by Materials Project with 10, and OMDB with one. Details on sample counts are in Section 3.2.
LLM4Mat-Bench provides the most comprehensive dataset compared to existing benchmarks, with the largest number of samples, properties, and tasks, including 60 regression and 5 classification tasks (see Table 2). It also offers more diverse material representations, incorporating chemical formulas, crystal structures, and crystal text descriptions. In contrast, MatBench (Dunn et al., 2020) and TextEdge (Rubungo et al., 2023) have fewer tasks and less representation diversity, with MatBench lacking crystal text descriptions and TextEdge missing material compositions and crystal structures.
2.3 Data Quality
Since Robocrystallographer generates crystal text descriptions in a deterministic manner following predefined and well-validated rules (Jain et al., 2013), these texts should faithfully describe the crystal structures used to generate them. Regarding the quality of labels, they are calculated from simulations and are usually considered noise-free. Properties data except those from JARVIS-QETB and hMOF are obtained from DFT, which is based on fundamental quantum mechanical equations. While DFT calculations can still be performed with different levels of approximations and fidelity, the DFT-calculated properties are usually considered to be highly reliable and are routinely used as noise-free ground truths for ML models in the materials science community.
3 Results
3.1 Experimental Details
We conducted about 845 experiments, evaluating the performance of five models and three material representations on each property for each data source. Consistent with standard practices in materials science, we evaluated performance separately for each data source rather than combining samples from different sources for the same property. This approach accounts for variations in techniques and settings used by different data sources, which can result in discrepancies, such as differing band gaps for the same material. Below, we will describe each material representation, model, and metric that we used to conduct our experiments.
3.1.1 Material Representations
LLM4Mat-Bench includes three distinct materials representations: Composition, CIF, and Description (see Table A). The primary goal of using these diverse representations is to identify which best enhances LLM performance in predicting material properties across different data sources.
Composition (Comp.) Material composition refers to the chemical formula of a material. Though it only provides stoichiometric information, studies have shown it can still be a reliable material representation for property prediction (Dunn et al., 2020; Tian et al., 2022). For LLMs, it offers the advantage of being a short sequence that usually fits within the model’s context window, making it efficient to train. To further optimize efficiency, we set the longest sequence of material compositions from each data source as the context window, rather than using the default 512 tokens for fine-tuning while the original length is kept during inference.
CIF We represent the materials structure using CIF files, the conventional way of representing the crystal structure in crystallography (Hall et al., 1991). CIFs are commonly used for GNN-based models, but some recent works have demonstrated that it can also work with LLMs (Antunes et al., 2023; Flam-Shepherd and Aspuru-Guzik, 2023; Gruver et al., 2023).
Description (Descr.) As we outlined in Section 2.1.3, we also use textual descriptions of crystal structures as representations for both atomic crystals and MOFs.
3.1.2 Models
We benchmarked different LLM-based models with various sizes, and a GNN-based baseline. Herein, We provide the details of each model.
CGCNN (Xie and Grossman, 2018) is employed as a GNN baseline which is widely used in the materials science community [^16] Although CGCNN is not state-of-the-art for some properties, it was faster compared to models like ALIGNN (Choudhary and DeCost, 2021) and DeeperGatGNN (Omee et al., 2022), making it suitable for our extensive experiments. We trained on LLM4Mat-Bench from scratch with optimal hyperparameters: 128 hidden dimensions, batch size of 256, three message passing layers, 1e-2 learning rate, 8.0 radius cutoff, 12 nearest neighbors, and 500 training epochs, though extending to 1000 epochs improved performance in some cases.
MatBERT (Walker et al., 2021) is a BERT-base model (Devlin et al., 2018) with 109 million parameters, pretrained on two million materials science articles. We fine-tuned MatBERT on LLM4Mat-Bench, following Rubungo et al. (2023), and achieved optimal performance with a 512-token input length, 64-sample batch size, 5e-5 learning rate, 0.5 dropout, and 100 epochs using the Adam optimizer and onecycle learning rate scheduler (Smith and Topin, 2019). Although training for 200 epochs improves performance, results are reported for 100 epochs due to computational constraints.
LLM-Prop is a model based on the encoder part of T5-small model (Raffel et al., 2020) introduced by Rubungo et al. (2023), with 35 million parameters, smaller than MatBERT. It predicts material properties from the textual descriptions of crystal structures. To adapt LLM-Prop on CIF, we employed xVal encoding (Golkar et al., 2023), where we parse an input sequence $x$ to extract numerical values into a list $x_{\text{num}}$, replace them with a [NUM] token to form $x_{\text{text}}$, and then embed $x_{\text{text}}$, followed by multiplying each [NUM] embedding by its corresponding value in $x_{\text{num}}$ to get $h_{\text{embed}}$ that we feed to the model. xVal encoding ensures that the quantitative value of each number is reflected in the input embedding while reducing the input length caused by the high volume of numerical values in CIF files, which extend the length of the input sequence after tokenization. We fine-tuned LLM-Prop on LLM4Mat-Bench and optimizing with a 1e-3 learning rate, 0.2 dropout, Adam optimizer, and onecycle learning rate scheduler for 100 epochs, with a 768-token input length, batch size of 64 for training, and 512 for inference. While Rubungo et al. (2023) recommended that training for 200 epochs and increasing the number of input tokens improves the performance, we could not replicate this due to computational constraints.
Table 3: Prompt template. <material representation type> denotes “chemical formula", “cif structure", or “structure description". <value> represents the input context (for example NaCl, etc.). <property name> denotes the name of the property (for example band gap, etc.). <predicted value> represents the property value generated by Llama 2 while <actual value_i> represents the ground truth of the EXAMPLE_i. FINAL PROMPT and RESPONSE denote the input prompt to Llama 2 and its generated output, respectively.
Prompt Type Template
- SYSTEM PROMPT:
<<SYS>>
You are a material scientist.
Look at the <material representation type> of the given crystalline material and predict its property.
The output must be in a json format. For example: {property_name:predicted_property_value}.
Answer as precise as possible and in as few words as possible.
<</SYS>>
INPUT PROMPT:
<material representation type>: <value>
property name: <property name>.
0-shot FINAL PROMPT: <s>[INST] + SYSTEM PROMPT + INPUT PROMPT + [/INST]
RESPONSE: <property name>:<predicted value>
5-shot EXAMPLE_1:                 …     EXAMPLE_5:
<material representation type>: <value_1>       <material representation type>: <value_5>
property name: <property name>.           property name: <property name>.
<property name>:<actual value_1>            <property name>:<actual value_5>
FINAL PROMPT: <s>[INST] + SYSTEM PROMPT + EXAMPLE_1 + … + EXAMPLE_5 + INPUT PROMPT + [/INST]
RESPONSE: <property name>:<predicted value>
Llama 2-7b-chat To assess the performance of conversational LLMs in materials property prediction, we tested Llama 2-7b-chat (7 billion parameters) using our designed zero-shot and five-shot prompts (see Table 3) without fine-tuning. For the CIF structure prompts, we removed "# generated using pymatgen" comment that is appended to each file. The maximum input length was set to 4000 tokens while the output length was set to 256, with a batch size of 256 samples, temperature of 0.8, and top-K sampling applied with $K=10$. The details of other models that we compared with Llama 2 can be found in Appendix C. For five-shot examples, we sampled from crystals with shorter structures and descriptions to reduce the context length. We also made sure the property values for those examples are diverse (for instance, they should not all have 0.0 eV as their bandgap values).
We trained all models using NVIDIA RTX A6000 GPUs. Training MatBERT with two GPUs on about 300K data points and 100 epochs took about four days while for LLM-Prop took about 2.5 days. For CGCNN, it took about 7 hours training time on one GPU for 500 epochs. With one GPU, Llama 2 took about a half day to generate the output of 40K samples with 256 tokens maximum length each. We report the test set results averaged over five runs for predictive models and three runs for generative models.
3.1.3 Evaluation Metrics
Following Choudhary and DeCost (2021), we evaluated regression tasks using the ratio between the mean absolute deviation (MAD) of the ground truth and the mean absolute error (MAE) of the predicted properties. The MAD:MAE ratio ensures an unbiased model comparison between different properties where the higher ratio the better. According to Choudhary and DeCost (2021), a good predictive model should have at least 5.0 ratio. MAD values represent the performance of a random guessing model predicting the average value for each data point. To provide a comprehensive performance comparison across datasets, we also reported the weighted average of MAD:MAE across all properties in each dataset (Wtd. Avg. (MAD:MAE), see Equation 1).
For classification tasks, we reported the area under the ROC curve (AUC) for each task and provided the weighted average across all properties (Wtd. Avg. AUC, see Equation 2).
[…]
$m$ denotes the number of regression properties in the dataset.
Table 4: The Wtd. Avg. (MAD/MAE) scores (the higher the better) for the regression tasks in the LLM4Mat-Bench are reported. Bolded results indicate the best model for each input format, while bolded results with blue background highlight the best model per each dataset. Inval. denotes cases where the Llama model failed to generate outputs with a property value or had fewer than 10 valid predictions.
Input Model Dataset
MP JARVIS-DFT GNoME hMOF Cantor HEA JARVIS-QETB OQMD QMOF SNUMAT OMDB
8 tasks 20 tasks 6 tasks 7 tasks 4 tasks 4 tasks 2 tasks 4 tasks 4 tasks 1 task
CIF CGCNN (baseline) 5.319 7.048 19.478 2.257 17.780 61.729 14.496 3.076 1.973 2.751
Comp. Llama 2-7b-chat:0S 0.389 Inval. 0.164 0.174 0.034 0.188 0.105 0.303 0.940 0.885
Llama 2-7b-chat:5S 0.627 0.704 0.499 0.655 0.867 1.047 1.160 0.932 1.157 1.009
MatBERT-109M 5.317 4.103 12.834 1.430 6.769 11.952 5.772 2.049 1.828 1.554
LLM-Prop-35M 4.394 2.912 15.599 1.479 8.400 59.443 6.020 1.958 1.509 1.507
CIF Llama 2-7b-chat:0S 0.392 0.216 6.746 0.214 0.022 0.278 0.028 0.119 0.682 0.159
Llama 2-7b-chat:5S Inval. Inval. Inval. Inval. Inval. 1.152 1.391 Inval. Inval. 0.930
MatBERT-109M 7.452 6.211 14.227 1.514 9.958 47.687 10.521 3.024 2.131 1.777
LLM-Prop-35M 8.554 6.756 16.032 1.623 15.728 97.919 11.041 3.076 1.829 1.777
Descr. Llama 2-7b-chat:0S 0.437 0.247 0.336 0.193 0.069 0.264 0.106 0.152 0.883 0.155
Llama 2-7b-chat:5S 0.635 0.703 0.470 0.653 0.820 0.980 1.230 0.946 1.040 1.001
MatBERT-109M 7.651 6.083 15.558 1.558 9.976 46.586 11.027 3.055 2.152 1.847
LLM-Prop-35M 9.116 7.204 16.224 1.706 15.926 93.001 9.995 3.016 1.950 1.656
Table 5: The Wtd. Avg. AUC scores (the higher the better) for the classification tasks in the LLM4Mat-Bench.
Input Model Dataset
MP SNUMAT
2 tasks 3 tasks
CIF CGCNN (baseline) 0.846 0.722
Comp. Llama 2-7b-chat:0S 0.491 Inval.
Llama 2-7b-chat:5S 0.507 0.466
MatBERT-109M 0.722 0.712
LLM-Prop-35M 0.691 0.716
CIF Llama 2-7b-chat:0S 0.501 0.489
Llama 2-7b-chat:5S 0.502 0.474
MatBERT-109M 0.750 0.717
LLM-Prop-35M 0.738 0.660
Descr. Llama 2-7b-chat:0S 0.500 Inval.
Llama 2-7b-chat:5S 0.502 0.568
MatBERT-109M 0.735 0.730
LLM-Prop-35M 0.742 0.735
Figure 1: The performance comparison across models for each material representation is presented. The left y-axis shows the log-normalized performance of each LLM-based model relative to the baseline (CGCNN), while the right y-axis (bar plots) displays the average subword tokens per sample for each dataset. Datasets on the x-axis are ordered by increasing average subword tokens. Results for Llama 2-7b-chat:0S and Llama 2-7b-chat:5S are missing in plots (a) and (b), respectively, due to invalid outputs. Higher values in the line plots indicate better performance.
Figure 2: The performance comparison across material representations for each LLM-based model is shown. The y-axis represents the log-normalized Weighted Average (MAD/MAE) score for each representation, while the x-axis displays randomly ordered datasets. In the (a) and (b) plots, some Composition and Structure performance results are missing due to invalid outputs. A higher y-axis value indicates better performance.
Figure 3: The performance comparison of different chat-based LLM versions is presented with results based on 5-shot prompts, averaged over three inference runs. Panels (a)–(c) and (d)–(f) show each model’s accuracy in predicting band gaps and stability in the MP dataset, respectively, while panels (g)–(i) and (j)–(l) depict the percentage of valid predictions for band gap and stability on the test set.
3.2 Discussion
Table 4 and 5, and Figure 1 and 2 show the main results. The detailed results on each dataset can be found in Appendix E. The main observations are as follows:
Small, task-specific predictive LLMs exhibit significantly better performance than larger, generative general-purpose LLMs. This performance disparity is evident across both regression (Table 4 and Figure 1) and classification tasks (Table 5) on all 10 datasets. Specifically, LLM-Prop and MatBERT outperform Llama 2-7b-chat:0S and Llama 2-7b-chat:5S by a substantial margin, despite being approximately 200 and 64 times smaller in size, respectively. In regression tasks, LLM-Prop achieves the highest accuracy on 8 out of 10 datasets, with MatBERT leading on the remaining 2 datasets. For classification tasks, both LLM-Prop and MatBERT deliver the best performance on 1 out of 2 datasets. LLM-Prop surpasses MatBERT by 1.8% on the SNUMAT dataset, whereas MatBERT outperforms LLM-Prop by 0.8% on the other dataset. As anticipated, a modest enhancement in average performance is observed across various datasets and input formats when the Llama 2-7b-chat model is evaluated using 5-shot prompts rather than 0-shot prompts. Determining the optimal number of examples required to achieve peak performance will be the focus of future work.
General-purpose generative LLMs hallucinate and often fail to generate valid property values. As shown in Table 4, Table 5, Figure 2, and Appendix E, Llama 2-7b-chat model produces invalid outputs on multiple tasks, where the expected property value is missing. This issue occurs less frequently when the input is a description or chemical formula, but more commonly when the input is a CIF file. One reason may be that descriptions and chemical formulas resemble natural language, which LLMs can more easily interpret compared to CIF files. This may also indicate that when the input modality during inference differs significantly from the modalities encountered during pretraining, fine-tuning is necessary to achieve reasonable performance. Another key observation is that Llama 2-7b-chat model often generates the same property value for different inputs (i.e. hallucinates), contributing to its poor performance across multiple tasks. These findings highlight the importance of caution when using general-purpose generative LLMs for materials property prediction and emphasize the need for fine-tuned, task-specific LLM-based models.
Representing materials with their textual descriptions improves the performance of LLM-based property predictors compared to other representations. We observe a significant performance improvement when the input is a description compared to when it is a CIF file or a chemical formula. One of the possible reasons for this might be that LLMs are more adept at learning from natural language data. On the other hand, although material compositions appear more natural to LLMs compared to CIF files, they lack sufficient structural information. This is likely why LLMs with CIF files as input significantly outperform those using chemical formulas.
More advanced, general-purpose generative LLMs do not necessarily yield better results in predicting material properties. In Figure 3, we compare the performance of Llama 2-7b-chat-hf model with advanced versions of Llama of comparable sizes when predicting material’s band gap and its stability. Similar comparisons are also conducted for the Mistral (Jiang et al., 2023) and Gemma (Team et al., 2024) models. The results indicate that, despite being trained on substantially larger and higher-quality datasets, more advanced versions of generative LLMs show limited improvements in performance and validity of predictions for material properties. For instance, Llama 3 and 3.1 8b models were trained on over 15 trillion tokens—around eight times more data than the 2 trillion tokens used for the Llama 2 7b models. This finding highlights the ongoing challenges of leveraging LLMs in material property prediction and underscores the need for further research to harness the potential of these robust models in this domain.
The performance on energetic properties is consistently better across all datasets compared to other properties. This is consistent with the trend observed in the community benchmarks such as MatBench and JARVIS-Leaderboard, where energetic properties are among those that can be most accurately predicted (Dunn et al., 2020; Choudhary et al., 2024). This is not surprising because energy is known to be relatively well predicted from e.g., compositions and atom coordination (bonding), which is inherently represented in GNNs and also presented in text descriptions.
Task-specific predictive LLM-based models excel with shorter textual descriptions, while CGCNN performs better on datasets with longer descriptions. While the focus on this work is on LLMs, a comparison with a simple but widely used GNN-based baseline suggests room for improvement in LLM-based property prediction. For regression tasks, LLM-Prop outperforms CGCNN on only 4 out of 10 datasets (MP, JARVIS-DFT, JARVIS-QETB, and SNUMAT), and MatBERT outperforms CGCNN on just 2 out of 10 datasets (MP and JARVIS-QETB). In contrast, CGCNN achieves the best performance on 5 out of 10 datasets (GNoME, hMOF, Cantor HEA, OQMD, and OMDB). Further analysis reveals that CGCNN tends to perform better than LLM-based models on datasets with relatively longer textual descriptions, while LLM-based models excel on datasets with shorter descriptions (see Table 1). The performance gain on shorter descriptions may stem from LLM-based models’ ability to leverage more context from compact text, while CGCNN consistently benefits from training on the entire crystal structure.
4 Conclusion
LLMs are increasingly being utilized in materials science, particularly for materials property prediction and discovery. However, the absence of standardized evaluation benchmarks has impeded progress in this field. We introduced LLM4Mat-Bench, a comprehensive benchmark dataset designed to evaluate LLMs for predicting properties of atomic and molecular crystals and MOFs. Our results demonstrate the limitations of general-purpose LLMs in this domain and underscore the necessity for task-specific predictive models and instruction-tuned LLMs tailored for materials property prediction. These findings emphasize the importance of using LLM4Mat-Bench to advance the development of more effective LLMs in materials science.
5 Limitations
Due to computational constraints and the number of experiments, we were unable to conduct thorough hyperparameter searches for each property and dataset. The reported settings were optimized on the MP dataset and then fixed for other datasets. For each model, we highlighted hyperparameter settings that may improve performance (see Section 3.1.2). Additionally, we could not include results from SOTA commercial LLMs such as GPT-4o [^17] https://openai.com/index/hello-gpt-4o/ or Claude 3.5 Sonnet [^18] https://www.anthropic.com/news/claude-3-5-sonnet due to budget constraints.
We also encountered issues with chat-based models, which sometimes failed to follow the output format, producing invalid or incomplete outputs. Extracting property values was therefore challenging. We believe further instruction-tuning chat-based models on the provided prompts could mitigate these issues.
Furthermore, we did not include comparisons with dataset-specific retrieval-augmented generation (RAG) models, such as the recently developed LLaMP (Chiang et al., 2024), a RAG-based model tailored for interaction with the MP dataset. Our work aims to provide a comprehensive benchmark and baseline results to advance the evaluation of LLM-based methods for materials property prediction. Future work should address these limitations.
Acknowledgments
Adji Bousso Dieng acknowledges support from the National Science Foundation, Office of Advanced Cyberinfrastructure (OAC) #2118201, and from the Schmidt Sciences AI2050 Early Career Fellowship.
Appendices
[…]
Appendix A Materials Representations
Table 6: LLM4Mat-Bench material representations of Sodium Chloride (NaCl).
Crystal Information File (CIF)
1# generated using pymatgen 2data_NaCl 3_symmetry_space_group_name_H-M P 1’ 4_cell_length_a 3.50219000 5_cell_length_b 3.50219000 6_cell_length_c 3.50219000 7_cell_angle_alpha 90.00000000 8_cell_angle_beta 90.00000000 9_cell_angle_gamma 90.00000000 10_symmetry_Int_Tables_number 1 11_chemical_formula_structural NaCl 12_chemical_formula_sum Na1 Cl1 13_cell_volume 42.95553287 14_cell_formula_units_Z 1 15loop_ 16 _symmetry_equiv_pos_site_id 17 _symmetry_equiv_pos_as_xyz 18 1 x, y, z 19loop_ 20 _atom_type_symbol 21 _atom_type_oxidation_number 22 Na+ 1.0 23 Cl- -1.0 24loop_ 25 _atom_site_type_symbol 26 _atom_site_label 27 _atom_site_symmetry_multiplicity 28 _atom_site_fract_x 29 _atom_site_fract_y 30 _atom_site_fract_z 31 _atom_site_occupancy 32 Na+ Na0 1 0.00000000 0.00000000 0.00000000 1 33 Cl- Cl1 1 0.50000000 0.50000000 0.50000000 1
Description
NaCl is Tetraauricupride structured and crystallizes in the cubic $P\overline{m}3m$ space group. Na1+ is bonded in a body-centered cubic geometry to eight equivalent Cl1- atoms. All Na-Cl bond lengths are 3.03 Å. Cl1- is bonded in a body-centered cubic geometry to eight equivalent Na1+ atoms.
Appendix B Statistics of All Properties in LLM4Mat-Bench
Table 7: Statistics of all datasets in LLM4Mat-Bench. It is important to note that we retain the naming convention of each property from the original data source with the intent to provide the distribution of properties in each dataset.
Property Task type # Samples/Data source
JARVIS-DFT Materials Project SNUMAT hMOF GNoME JARVIS-QETB Cantor HEA QMOF OQMD OMDB
Bandgap Regression - 145,302 - - 288,209 - - 16,340 1,007,324 12,500
Bandgap (OPT) Regression 75,965 - - - - - - - - -
Bandgap (MBJ) Regression 19,800 - - - - - - - - -
Bandgap GGA Regression - - 10,481 - - - - - - -
Bandgap HSE Regression - - 10,481 - - - - - - -
Bandgap GGA Optical Regression - - 10,481 - - - - - - -
Bandgap HSE Optical Regression - - 10,481 - - - - - - -
Indirect Bandgap Regression - - - - - 829,576 - - - -
Formation Energy Per Atom (FEPA) Regression 75,965 145,262 - - 384,871 829,576 84,024 - 1,008,266 -
Energy Per Atom (EPA) Regression - 145,262 - - - 829,576 84,024 - - -
Decomposition Energy Per Atom (DEPA) Regression - - - - 384,871 - - - - -
Energy Above Hull (Ehull) Regression 75,965 145,262 - - - - 84,024 - - -
Total Energy Regression 75,965 - - - 384,871 829,576 - 16,340 - -
Efermi Regression - 145,262 - - - - - - - -
Exfoliation Energy Regression 812 - - - - - - - - -
Bulk Modulus (Kv) Regression 23,823 - - - - - - - - -
Shear Modulus (Gv) Regression 23,823 - - - - - - - - -
SLME Regression 9,765 - - - - - - - - -
Spillage Regression 11,377 - - - - - - - - -
$\epsilon_{x}$ (OPT) Regression 52,158 - - - - - - - - -
$\epsilon$ (DFPT) Regression 4,704 - - - - - - - - -
Max. piezoelectri c strain coeff (dij) Regression 3,347 - - - - - - - - -
Max. piezo. stress coeff (eij) Regression 4,797 - - - - - - - - -
Max. EFG Regression 11,871 - - - - - - - - -
Avg. $m_{e}$ Regression 17,643 - - - - - - - - -
Is Stable Classification - 145,262 - - - - - - - -
Is Gap Direct Classification - 145,262 - - - - - - - -
n-Seedbeck Regression 23,211 - - - - - - - - -
n-PF Regression 23,211 - - - - - - - - -
p-Seedbeck Regression 23,211 - - - - - - - - -
p-PF Regression 23,211 - - - - - - - - -
Density Regression - 145,262 - - 384,871 - - - - -
Density Atomic Regression - 145,262 - - - - - - - -
Volume Regression - 145,262 - - 384,871 - - - - -
Volume Per Atom (VPA) Regression - - - - - - 84,024 - - -
Is Direct Classification - - 10,481 - - - - - - -
Is Direct HSE Classification - - 10,481 - - - - - - -
SOC Classification - - 10,481 - - - - - - -
LCD Regression - - - 133,524 - - - 16,340 - -
PLD Regression - - - 133,524 - - - 16,340 - -
Max CO2 Regression - - - 133,524 - - - - - -
Min CO2 Regression - - - 133,524 - - - - - -
Void Fraction Regression - - - 133,524 - - - - - -
Surface Area m2g Regression - - - 133,524 - - - - - -
Surface Area m2cm3 Regression - - - 133,524 - - - - - -
Appendix C Chat-like Model Inference Details
Table 8: Hyperparameters used during inference. Temp. represents temperature.
Model Type Model Name Input Length Output Length Batch Size Temp. Top_K
Llama Family Llama 2-7b-chat-hf 4000 256 256 0.8 10
Llama 3-8b-Instruct 8000 256 256 0.8 10
Llama 3.1-8b-Instruct 98000 256 128 0.8 10
Llama 3.2-3b-Instruct 47000 256 128 0.8 10
Gemma Family Gemma 7b-it 4000 256 256 0.8 10
Gemma 1.1-7b-it 4000 256 256 0.8 10
Gemma 2-9b-it 3000 256 256 0.8 10
Mistral Family Mistral 7b-Instruct-v0.1 20000 256 256 0.8 10
Mistral 7b-Instruct-v0.2 20000 256 256 0.8 10
Mistral 7b-Instruct-v0.3 20000 256 256 0.8 10
Appendix D Prompt Templates
Figure 4: Prompt templates when the input is a chemical formula.
Figure 5: Prompt templates when the input is a CIF file.
Figure 6: Prompt templates when the input is a crystal structure description.
Appendix E Result Details for Each Dataset
E.1 Detailed MAD:MAE Results
Table 9: Results for MP dataset. The performance on regression tasks is evaluated in terms of MAD:MAE ratio (the higher the better) while that of classification tasks (Is Stable and Is Gab Direct) is evaluated in terms of AUC score. FEPA: Formation Energy Per Atom, EPA: Energy Per Atom.
Input Model FEPA Bandgap EPA Ehull Efermi Density Density Atomic Volume Is Stable Is Gab Direct
145.2K 145.3K 145.2K 145.2K 145.2K 145.2K 145.2K 145.2K 145.2K 145.2K
CIF CGCNN (baseline) 8.151 3.255 7.224 3.874 3.689 8.773 5.888 1.703 0.882 0.810
Comp. Llama 2-7b-chat:0S 0.008 0.623 0.009 0.001 0.003 0.967 0.754 0.747 0.500 0.482
Llama 2-7b-chat:5S 0.33 1.217 0.239 0.132 0.706 0.899 0.724 0.771 0.502 0.512
MatBERT-109M 8.151 2.971 9.32 2.583 3.527 7.626 5.26 3.099 0.764 0.681
LLM-Prop-35M 7.482 2.345 7.437 2.006 3.159 6.682 3.523 2.521 0.746 0.636
CIF Llama 2-7b-chat:0S 0.032 0.135 0.022 0.001 0.015 0.97 0.549 1.41 0.503 0.499
Llama 2-7b-chat:5S Inval. 1.111 0.289 Inval. 0.685 0.98 0.99 0.926 0.498 0.506
MatBERT-109M 11.017 3.423 13.244 3.808 4.435 10.426 6.686 6.58 0.790 0.710
LLM-Prop-35M 14.322 3.758 17.354 2.182 4.515 13.834 4.913 7.556 0.776 0.700
Descr. Llama 2-7b-chat:0S 0.019 0.633 0.023 0.001 0.008 1.31 0.693 0.807 0.500 0.500
Llama 2-7b-chat:5S 0.394 1.061 0.297 0.247 0.684 0.916 0.782 0.704 0.500 0.504
MatBERT-109M 11.935 3.524 13.851 4.085 4.323 9.9 6.899 6.693 0.794 0.713
LLM-Prop-35M 15.913 3.931 18.412 2.74 4.598 14.388 4.063 8.888 0.794 0.690
Table 10: Results for JARVIS-DFT. he performance on regression tasks is evaluated in terms of MAD:MAE ratio (the higher the better). FEPA: Formation Energy Per Atom, Tot. En.: Total Energy, Exf. En.: Exfoliation Energy.
Input Model FEPA Bandgap (OPT) Tot. En. Ehull Bandgap (MBJ) Kv Gv SLME Spillage $\epsilon_{x}$ (OPT)
75.9K 75.9K 75.9K 75.9K 19.8K 23.8K 23.8K 9.7K 11.3K 18.2K
CIF CGCNN (baseline) 13.615 4.797 22.906 1.573 4.497 3.715 2.337 1.862 1.271 2.425
Comp. Llama 2-7b-chat:0S 0.021 0.011 0.02 0.005 0.92 0.428 0.374 0.148 Inval. 0.18
Llama 2-7b-chat:5S 0.886 0.011 0.02 1.292 0.979 0.88 0.992 0.456 0.85 1.148
MatBERT-109M 6.808 4.083 9.21 2.786 3.755 2.906 1.928 1.801 1.243 2.017
LLM-Prop-35M 4.765 2.621 5.936 2.073 2.922 2.162 1.654 1.575 1.14 1.734
CIF Llama 2-7b-chat:0S 0.023 0.011 0.02 0.002 0.193 0.278 0.358 0.186 0.702 0.781
Llama 2-7b-chat:5S 0.859 Inval. Inval. 1.173 1.054 0.874 0.91 0.486 0.916 1.253
MatBERT-109M 10.211 5.483 15.673 4.862 5.344 4.283 2.6 2.208 1.444 2.408
LLM-Prop-35M 12.996 3.331 22.058 2.648 4.93 4.121 2.409 2.175 1.37 2.135
Descr. Llama 2-7b-chat:0S 0.007 0.011 0.02 0.004 0.94 0.498 0.382 0.07 0.135 0.647
Llama 2-7b-chat:5S 0.845 0.011 0.02 1.273 1.033 0.87 0.969 0.461 0.857 1.201
MatBERT-109M 10.211 5.33 15.141 4.691 5.01 4.252 2.623 2.178 1.452 2.384
LLM-Prop-35M 12.614 3.427 23.509 4.532 4.983 4.128 2.419 2.061 1.307 2.334
$\epsilon$ (DFPT) Max. Piezo. (dij) Max. Piezo. (eij) Max. EFG Exf. En. Avg. $m_{e}$ n-Seebeck n-PF p-Seebeck p-PF
4.7K 3.3K 4.7K 11.8K 0.8K 17.6K 23.2K 23.2K 23.2K 23.2K
CIF CGCNN (baseline) 1.12 0.418 1.291 1.787 0.842 1.796 2.23 1.573 3.963 1.59
Comp. Llama 2-7b-chat:0S 0.012 0.121 0.001 0.141 0.384 0.028 0.874 0.801 0.971 0.874
Llama 2-7b-chat:5S 1.416 1.289 1.305 0.765 0.512 0.535 1.008 1.04 0.93 0.568
MatBERT-109M 1.533 1.464 1.426 1.658 1.124 2.093 1.908 1.318 2.752 1.356
LLM-Prop-35M 1.454 1.447 1.573 1.38 1.042 1.658 1.725 1.145 2.233 1.285
CIF Llama 2-7b-chat:0S 0.033 0.104 0.001 0.246 0.411 0.041 0.429 0.766 0.83 0.826
Llama 2-7b-chat:5S Inval. Inval. Inval. 0.796 0.51 Inval. 1.039 1.396 Inval. Inval.
MatBERT-109M 1.509 1.758 2.405 2.143 1.374 2.45 2.268 1.446 3.337 1.476
LLM-Prop-35M 1.578 2.103 2.405 1.936 1.044 1.796 1.955 1.332 2.503 1.399
Descr. Llama 2-7b-chat:0S 0.08 0.266 0.001 0.138 0.285 0.019 0.769 0.793 0.825 0.829
Llama 2-7b-chat:5S 1.649 1.174 1.152 0.806 0.661 0.523 1.098 1.024 0.948 0.563
MatBERT-109M 1.534 1.807 2.556 2.081 1.36 2.597 2.241 1.432 3.26 1.565
LLM-Prop-35M 1.64 2.116 2.315 1.978 1.168 1.858 2.154 1.364 2.61 1.407
Table 11: Results for SNUMAT. The performance on regression tasks is evaluated in terms of MAD:MAE ratio (the higher the better) while that of classification tasks (Is Direct, Is Direct HSE, and SOC) is evaluated in terms of AUC score.
Input Model Bandgap GGA Bandgap HSE Bandgap GGA Optical Bandgap HSE Optical Is Direct Is Direct HSE SOC
10.3K 10.3K 10.3K 10.3K 10.3K 10.3K 10.3K
CIF CGCNN (baseline) 2.075 2.257 1.727 1.835 0.691 0.675 0.800
Comp. Llama 2-7b-chat:0S 0.797 0.948 1.156 0.859 0.503 0.484 Inval.
Llama 2-7b-chat:5S 1.267 1.327 0.862 1.174 0.475 0.468 0.455
MatBERT-109M 1.899 1.975 1.646 1.793 0.671 0.645 0.820
LLM-Prop-35M 1.533 1.621 1.392 1.491 0.647 0.624 0.829
CIF Llama 2-7b-chat:0S 0.346 0.454 1.09 0.838 0.479 0.488 0.500
Llama 2-7b-chat:5S Inval. Inval. Inval. Inval. 0.494 0.500 0.427
MatBERT-109M 2.28 2.472 1.885 1.889 0.677 0.650 0.823
LLM-Prop-35M 1.23 2.401 1.786 1.9 0.661 0.664 0.656
Descr. Llama 2-7b-chat:0S 0.802 0.941 1.013 0.779 0.499 0.509 Inval.
Llama 2-7b-chat:5S 0.774 1.315 0.901 1.172 0.594 0.623 0.486
MatBERT-109M 2.298 2.433 1.901 1.978 0.683 0.645 0.862
LLM-Prop-35M 2.251 2.142 1.84 1.569 0.681 0.657 0.866
Table 12: Results for GNoME. The performance on regression tasks is evaluated in terms of MAD:MAE ratio (the higher the better). FEPA: Formation Energy Per Atom, DEPA: Decomposition Energy Per Atom, Tot. En.: Total Energy.
Input Model FEPA Bandgap DEPA Tot. En. Volume Density
376.2K 282.7K 376.2K 282.7K 282.7K 282.7K
CIF CGCNN (baseline) 34.57 8.549 2.787 7.443 7.967 56.077
Comp. Llama 2-7b-chat:0S 0.002 0.177 0.0 0.088 0.455 0.368
Llama 2-7b-chat:5S 0.194 0.086 0.255 0.765 1.006 0.865
MatBERT-109M 30.248 4.692 2.787 8.57 13.157 15.145
LLM-Prop-35M 25.472 3.735 1.858 21.624 16.556 25.615
CIF Llama 2-7b-chat:0S 0.003 0.045 0.0 0.706 43.331 0.794
Llama 2-7b-chat:5S Inval. 0.087 Inval. Inval. 1.029 0.878
MatBERT-109M 24.199 9.16 3.716 15.309 16.691 16.467
LLM-Prop-35M 28.469 3.926 3.344 17.837 17.082 25.615
Descr. Llama 2-7b-chat:0S 0.002 0.114 0.0 0.661 0.654 0.805
Llama 2-7b-chat:5S 0.192 0.086 0.106 0.75 1.006 0.891
MatBERT-109M 30.248 5.829 3.716 18.205 17.824 16.599
LLM-Prop-35M 28.469 5.27 3.716 17.02 17.02 25.936
Table 13: Results for hMOF. The performance on regression tasks is evaluated in terms of MAD:MAE ratio (the higher the better).
Input Model Max CO2 Min CO2 LCD PLD Void Fraction Surface Area m2g Surface Area m2cm3
132.7K 132.7K 132.7K 132.7K 132.7K 132.7K 132.7K
CIF CGCNN (baseline) 1.719 1.617 1.989 1.757 2.912 3.765 2.039
Comp. Llama 2-7b-chat:0S 0.011 0.002 0.009 0.008 0.5 0.454 0.233
Llama 2-7b-chat:5S 0.679 0.058 0.949 1.026 0.945 0.567 0.366
MatBERT-109M 1.335 1.41 1.435 1.378 1.57 1.517 1.367
LLM-Prop-35M 1.41 1.392 1.432 1.468 1.672 1.657 1.321
CIF Llama 2-7b-chat:0S 0.017 0.003 0.016 0.011 0.549 0.54 0.359
Llama 2-7b-chat:5S Inval. Inval. 0.951 1.067 Inval. Inval. Inval.
MatBERT-109M 1.421 1.428 1.544 1.482 1.641 1.622 1.461
LLM-Prop-35M 1.564 1.41 1.753 1.435 1.9 1.926 1.374
Descr. Llama 2-7b-chat:0S 0.129 0.014 0.026 0.006 0.382 0.497 0.299
Llama 2-7b-chat:5S 0.684 0.058 0.955 1.006 0.931 0.571 0.37
MatBERT-109M 1.438 1.466 1.602 1.511 1.719 1.697 1.475
LLM-Prop-35M 1.659 1.486 1.623 1.789 1.736 2.144 1.508
Table 14: Results for Cantor HEA. The performance on regression tasks is evaluated in terms of MAD:MAE ratio (the higher the better). FEPA: Formation Energy Per Atom, EPA:Energy Per Atom, VPA:Volume Per Atom.
Input Model FEPA EPA Ehull VPA
84.0K 84.0K 84.0K 84.0K
CIF CGCNN (baseline) 9.036 49.521 9.697 2.869
Comp. Llama 2-7b-chat:0S 0.005 0.098 0.003 0.031
Llama 2-7b-chat:5S 0.896 0.658 0.928 0.986
MatBERT-109M 3.286 16.17 5.134 2.489
LLM-Prop-35M 3.286 22.638 5.134 2.543
CIF Llama 2-7b-chat:0S 0.001 0.084 0.0 0.004
Llama 2-7b-chat:5S Inval. Inval. Inval. Inval.
MatBERT-109M 7.229 17.607 9.187 5.809
LLM-Prop-35M 8.341 36.015 11.636 6.919
Descr. Llama 2-7b-chat:0S 0.001 0.101 0.164 0.011
Llama 2-7b-chat:5S 0.797 0.615 0.938 0.93
MatBERT-109M 7.229 17.607 9.187 5.881
LLM-Prop-35M 8.341 36.015 11.636 7.713
Table 15: Results for QMOF. The performance on regression tasks is evaluated in terms of MAD:MAE ratio (the higher the better). Tot. En.: Total Energy.
Input Model Bandgap Tot. En. LCD PLD
7.6K 7.6K 7.6K 7.6K
CIF CGCNN (baseline) 2.431 1.489 4.068 4.317
Comp. Llama 2-7b-chat:0S 0.901 0.26 0.045 0.009
Llama 2-7b-chat:5S 0.648 0.754 1.241 1.086
MatBERT-109M 1.823 1.695 2.329 2.349
LLM-Prop-35M 1.759 1.621 2.293 2.157
CIF Llama 2-7b-chat:0S 0.201 0.244 0.02 0.011
Llama 2-7b-chat:5S Inval. Inval. Inval. Inval.
MatBERT-109M 1.994 4.378 2.908 2.818
LLM-Prop-35M 2.166 4.323 2.947 2.87
Descr. Llama 2-7b-chat:0S 0.358 0.217 0.025 0.006
Llama 2-7b-chat:5S 0.777 0.713 1.125 1.17
MatBERT-109M 2.166 4.133 2.981 2.941
LLM-Prop-35M 2.091 4.312 2.831 2.829
Table 16: Results for JARVIS-QETB. The performance on regression tasks is evaluated in terms of MAD:MAE ratio (the higher the better). FEPA: Formation Energy Per Atom, EPA:Energy Per Atom, Tot. En.: Total Energy, Ind. Bandgap: Indirect Bandgap.
Input Model FEPA EPA Tot. En. Ind. Bandgap
623.9K 623.9K 623.9K 623.9K
CIF CGCNN (baseline) 1.964 228.201 11.218 5.534
Comp. Llama 2-7b-chat:0S 0.003 0.369 0.172 0.21
Llama 2-7b-chat:5S 0.812 1.037 1.032 1.306
MatBERT-109M 1.431 37.979 8.19 0.21
LLM-Prop-35M 2.846 211.757 21.309 1.861
CIF Llama 2-7b-chat:0S 0.003 0.412 0.656 0.04
Llama 2-7b-chat:5S 0.8 1.024 1.076 1.71
MatBERT-109M 24.72 135.156 26.094 4.779
LLM-Prop-35M 23.346 318.291 48.192 1.845
Descr. Llama 2-7b-chat:0S 0.003 0.408 0.484 0.16
Llama 2-7b-chat:5S 0.85 1.015 1.035 1.021
MatBERT-109M 26.265 122.884 29.409 7.788
LLM-Prop-35M 22.513 312.218 35.43 1.845
Table 17: Results for OQMD. The performance on regression tasks is evaluated in terms of MAD:MAE ratio (the higher the better). FEPA: Formation Energy Per Atom.
Input Model FEPA Bandgap
963.5K 963.5K
CIF CGCNN (baseline) 22.291 6.701
Comp. Llama 2-7b-chat:0S 0.019 0.192
Llama 2-7b-chat:5S 1.013 1.306
MatBERT-109M 7.662 3.883
LLM-Prop-35M 9.195 2.845
CIF Llama 2-7b-chat:0S 0.009 0.047
Llama 2-7b-chat:5S 1.051 1.731
MatBERT-109M 13.879 7.163
LLM-Prop-35M 18.861 3.22
Descr. Llama 2-7b-chat:0S 0.025 0.187
Llama 2-7b-chat:5S 0.991 1.468
MatBERT-109M 15.012 7.041
LLM-Prop-35M 16.346 3.644
Table 18: Results for OMDB. The performance on regression tasks is evaluated in terms of MAD:MAE ratio (the higher the better).
Input Model Bandgap
12.1K
CIF CGCNN (baseline) 2.751
Comp. Llama 2-7b-chat:0S 0.886
Llama 2-7b-chat:5S 1.009
MatBERT-109M 1.554
LLM-Prop-35M 1.507
CIF Llama 2-7b-chat:0S 0.159
Llama 2-7b-chat:5S 0.930
MatBERT-109M 1.777
LLM-Prop-35M 1.777
Descr. Llama 2-7b-chat:0S 0.155
Llama 2-7b-chat:5S 1.002
MatBERT-109M 1.847
LLM-Prop-35M 1.656
E.2 Detailed MAE Results
Table 19: Results for Materials Project (MP). The $\downarrow$ and $\uparrow$ on the right of each property denote the MAE error and AUC score, respectively. The Dummy model represents the MAD score that is calculated on the ground truth of each property.
Input Model FEPA $\downarrow$ Bandgap $\downarrow$ EPA $\downarrow$ Ehull $\downarrow$ Efermi $\downarrow$ Density $\downarrow$ Density atomic $\downarrow$ Volume $\downarrow$ Is stable $\uparrow$ Is gab direct $\uparrow$ Wtd. Avg. (MAD/MAE) $\uparrow$ Wtd. Avg. AUC $\uparrow$
- Dummy (MAD) 1.003 1.191 1.51 0.225 2.239 2.158 11.693 342.606 - - - -
Structure CGCNN (baseline) 0.123 0.366 0.209 0.058 0.607 0.246 1.986 201.186 0.882 0.810 5.319 0.846
Composition Llama 2-7b-chat:0S 128.007 1.913 163.899 195.908 673.939 2.231 15.503 458.456 0.500 0.482 0.389 0.491
Llama 2-7b-chat:5S 3.034 0.979 6.330 1.706 3.172 2.400 16.150 444.378 0.502 0.512 0.627 0.507
MatBERT-109M 0.123 0.401 0.162 0.087 0.635 0.283 2.223 110.556 0.764 0.681 5.316 0.722
LLM-Prop-35M 0.134 0.508 0.203 0.112 0.709 0.323 3.319 135.924 0.746 0.636 4.364 0.691
Structure Llama 2-7b-chat:0S 31.515 8.819 69.840 213.547 148.541 2.224 21.296 243.050 0.503 0.499 0.392 0.501
Llama 2-7b-chat:5S Inval. 1.072 5.219 Inval. 3.268 2.203 11.806 370.160 0.498 0.506 Inval. 0.502
MatBERT-109M 0.091 0.348 0.114 0.059 0.505 0.207 1.749 52.068 0.790 0.710 7.452 0.750
LLM-Prop-35M 0.070 0.317 0.087 0.103 0.496 0.156 2.380 45.340 0.776 0.700 8.554 0.738
Description Llama 2-7b-chat:0S 53.083 1.881 66.971 313.255 288.275 1.647 16.874 424.737 0.500 0.500 0.437 0.500
Llama 2-7b-chat:5S 2.544 1.123 5.087 0.910 3.274 2.356 14.961 486.650 0.500 0.504 0.635 0.502
MatBERT-109M 0.084 0.338 0.109 0.055 0.518 0.218 1.695 51.189 0.794 0.713 7.651 0.735
LLM-Prop-35M 0.063 0.303 0.082 0.082 0.487 0.150 2.878 38.548 0.794 0.690 9.116 0.742
Table 20: Results for JARVIS-DFT. The $\downarrow$ on the right of each property denotes the MAE error.
Input Model FEPA $\downarrow$ Bandgap (OPT) $\downarrow$ Total energy $\downarrow$ Ehull $\downarrow$ Bandgap (MBJ) $\downarrow$ Kv $\downarrow$ Gv $\downarrow$ SLME $\downarrow$ Spillage $\downarrow$ $\epsilon_{x}$ (OPT) $\downarrow$ Wtd. Avg. (MAD/MAE) $\uparrow$
- Dummy (MAD) 0.858 0.959 1.787 0.267 1.844 53.762 27.455 10.538 0.508 65.886 -
Structure CGCNN (baseline) 0.063 0.200 0.078 0.170 0.410 14.470 11.750 5.660 0.400 27.170 7.048
Composition Llama 2-7b-chat:0S 40.722 86.989 91.114 53.109 2.005 125.545 73.335 71.208 Inval. 365.525 Inval.
Llama 2-7b-chat:5S 0.968 87.324 90.542 0.207 1.883 61.121 27.672 23.093 0.598 57.394 0.704
MatBERT-109M 0.126 0.235 0.194 0.096 0.491 18.498 14.241 5.851 0.409 32.661 4.103
LLM-Prop-35M 0.180 0.366 0.301 0.129 0.631 24.868 16.603 6.689 0.446 37.993 2.912
Structure Llama 2-7b-chat:0S 37.893 86.424 90.286 135.879 9.532 193.714 76.784 56.666 0.724 84.332 0.216
Llama 2-7b-chat:5S 0.999 Inval. Inval. 0.228 1.749 61.509 30.156 21.678 0.555 52.576 Inval.
MatBERT-109M 0.084 0.175 0.114 0.055 0.345 12.553 10.560 4.772 0.352 27.364 6.211
LLM-Prop-35M 0.066 0.288 0.081 0.101 0.374 13.046 11.399 4.845 0.371 30.867 6.756
Description Llama 2-7b-chat:0S 117.691 87.327 90.615 72.709 1.962 108.062 71.895 150.412 3.753 101.906 0.247
Llama 2-7b-chat:5S 1.015 87.323 89.595 0.210 1.784 61.783 28.345 22.857 0.593 54.850 0.703
MatBERT-109M 0.084 0.180 0.118 0.057 0.368 12.645 10.466 4.839 0.350 27.633 6.083
LLM-Prop-35M 0.068 0.280 0.076 0.059 0.370 13.023 11.352 5.113 0.389 28.231 7.204
$\epsilon$ (DFPT) $\downarrow$ Max. piezo. (dij) $\downarrow$ Max. piezo. (eij) $\downarrow$ Max. EFG $\downarrow$ Exfoliation energy $\downarrow$ avg. $m_{e}$downarrow n-Seebeck $\downarrow$ n-PF $\downarrow$ p-Seebeck $\downarrow$ p-PF $\downarrow$
- Dummy (MAD) 43.451 14.525 0.245 44.129 42.08 0.216 111.33 695.882 168.11 699.799 -
Structure CGCNN (baseline) 38.780 34.710 0.190 24.700 50.000 0.120 49.920 442.300 42.420 440.26 7.048
Composition Llama 2-7b-chat:0S 3626.217 120.287 228.849 312.490 109.464 7.708 127.367 868.257 173.183 800.924 Inval.
Llama 2-7b-chat:5S 30.680 11.272 0.188 57.654 82.128 0.403 110.485 669.338 180.679 1231.228 0.704
MatBERT-109M 28.340 9.920 0.172 26.621 37.445 0.103 58.342 528.070 61.085 516.190 4.103
LLM-Prop-35M 29.881 10.037 0.156 31.966 40.385 0.130 64.538 608.007 75.268 544.737 2.912
Structure Llama 2-7b-chat:0S 1317.680 139.160 196.532 179.067 102.276 5.254 259.753 908.067 202.459 847.011 0.216
Llama 2-7b-chat:5S Inval. Inval. Inval. 55.417 82.580 Inval. 107.189 498.324 Inval. Inval. Inval.
MatBERT-109M 28.789 8.262 0.102 20.593 30.623 0.313 49.096 481.132 50.382 474.278 6.211
LLM-Prop-35M 27.529 6.908 0.102 22.791 40.324 0.120 56.953 522.337 67.153 500.348 6.756
Description Llama 2-7b-chat:0S 543.798 54.605 273.968 320.801 147.706 11.616 144.830 877.009 203.677 843.939 0.247
Llama 2-7b-chat:5S 26.354 12.374 0.213 54.739 63.690 0.412 101.373 679.530 177.250 1242.988 0.703
MatBERT-109M 28.329 8.036 0.096 21.202 30.948 0.083 49.674 486.017 51.570 447.099 6.083
LLM-Prop-35M 26.491 6.863 0.106 22.308 36.031 0.116 51.686 510.280 64.402 497.515 7.204
Table 21: Results for SNUMAT. The $\downarrow$ and $\uparrow$ on the right of each property denote the MAE error and AUC score, respectively.
Input Model Bandgap GGA $\downarrow$ Bandgap HSE $\downarrow$ Bandgap GGA optical $\downarrow$ Bandgap HSE optical $\downarrow$ Is direct $\uparrow$ Is direct HSE $\uparrow$ SOC $\uparrow$ Wtd. Avg. (MAD/MAE) $\uparrow$ Wtd. Avg. AUC $\uparrow$
- Dummy (MAD) 0.875 1.092 1.154 1.343 - - - - -
Structure CGCNN (baseline) 0.422 0.484 0.668 0.732 0.691 0.675 0.800 1.973 0.722
Composition Llama 2-7b-chat:0S 1.099 1.152 0.998 1.564 0.503 0.484 Inval. 0.940 Inval.
Llama 2-7b-chat:5S 0.691 0.823 1.339 1.144 0.475 0.468 0.455 1.157 0.466
MatBERT-109M 0.461 0.553 0.701 0.749 0.671 0.645 0.820 1.828 0.712
LLM-Prop-35M 0.571 0.674 0.829 0.901 0.647 0.624 0.829 1.509 0.716
Structure Llama 2-7b-chat:0S 2.531 2.405 1.058 1.603 0.479 0.488 0.500 0.682 0.489
Llama 2-7b-chat:5S Inval. Inval. Inval. Inval. 0.494 0.500 0.427 Inval. 0.474
MatBERT-109M 0.384 0.442 0.612 0.711 0.677 0.650 0.823 2.131 0.717
LLM-Prop-35M 0.712 0.455 0.646 0.707 0.661 0.664 0.656 1.829 0.660
Description Llama 2-7b-chat:0S 1.092 1.161 1.139 1.725 0.499 0.509 Inval. 0.883 Inval.
Llama 2-7b-chat:5S 1.131 0.831 1.281 1.146 0.594 0.623 0.486 1.040 0.568
MatBERT-109M 0.381 0.449 0.607 0.679 0.683 0.645 0.862 2.152 0.730
LLM-Prop-35M 0.389 0.510 0.627 0.856 0.681 0.657 0.866 1.950 0.735
Table 22: Results for GNoME. The $\downarrow$ on the right of each property denotes the MAE error.
Input Model FEPA $\downarrow$ Bandgap $\downarrow$ DEPA $\downarrow$ Total energy $\downarrow$ Volume $\downarrow$ Density $\downarrow$ Wtd. Avg. (MAD/MAE) $\uparrow$
- Dummy (MAD) 0.484 0.385 0.033 62.752 193.386 2.075 -
Structure CGCNN (baseline) 0.014 0.045 0.012 8.431 24.274 0.037 19.478
Composition Llama 2-7b-chat:0S 251.420 2.170 233.95 713.047 425.052 5.633 0.164
Llama 2-7b-chat:5S 2.494 4.471 0.131 81.987 192.289 2.398 0.499
MatBERT-109M 0.016 0.082 0.012 7.322 14.698 0.137 12.834
LLM-Prop-35M 0.019 0.103 0.018 2.902 11.681 0.081 15.599
Structure Llama 2-7b-chat:0S 150.755 8.481 77.790 88.822 4.463 2.612 6.746
Llama 2-7b-chat:5S Inval. 4.418 Inval. Inval. 188.008 2.364 Inval.
MatBERT-109M 0.020 0.042 0.009 4.099 11.586 0.126 14.227
LLM-Prop-35M 0.017 0.098 0.010 3.518 11.321 0.081 16.032
Description Llama 2-7b-chat:0S 259.528 3.360 311.944 94.968 295.904 2.576 0.336
Llama 2-7b-chat:5S 2.524 4.457 0.315 83.714 192.259 2.328 0.470
MatBERT-109M 0.016 0.066 0.009 3.447 10.850 0.125 15.558
LLM-Prop-35M 0.017 0.073 0.009 3.687 11.362 0.080 16.224
Table 23: Results for hMOF. The $\downarrow$ on the right of each property denotes the MAE error.
Input Model Max CO2 $\downarrow$ Min CO2 $\downarrow$ LCD $\downarrow$ PLD $\downarrow$ Void fraction $\downarrow$ Surface area m2g $\downarrow$ Surface area m2cm3 $\downarrow$ Wtd. Avg. (MAD/MAE) $\uparrow$
- Dummy (MAD) 2.245 0.11 3.506 3.59 0.181 1469.493 604.098 -
Structure CGCNN (baseline) 1.306 0.068 1.763 2.043 0.062 390.269 296.248 2.257
Composition Llama 2-7b-chat:0S 199.955 60.032 381.040 447.976 0.361 3239.754 2593.456 0.174
Llama 2-7b-chat:5S 3.305 1.902 3.696 3.500 0.191 2592.529 1651.846 0.655
MatBERT-109M 1.682 0.078 2.443 2.606 0.115 968.706 441.879 1.430
LLM-Prop-35M 1.592 0.079 2.449 2.445 0.108 886.750 457.435 1.479
Structure Llama 2-7b-chat:0S 128.932 32.271 221.708 327.567 0.329 2721.352 1682.345 0.214
Llama 2-7b-chat:5S Inval. Inval. 3.688 3.365 Inval. Inval. Inval. Inval.
MatBERT-109M 1.580 0.077 2.271 2.423 0.110 906.251 413.508 1.514
LLM-Prop-35M 1.435 0.078 2.000 2.501 0.095 763.111 439.739 1.623
Description Llama 2-7b-chat:0S 17.457 7.884 134.936 572.296 0.472 2956.530 2019.484 0.193
Llama 2-7b-chat:5S 3.281 1.891 3.671 3.568 0.194 2571.702 1634.674 0.653
MatBERT-109M 1.561 0.075 2.189 2.376 0.105 865.712 409.503 1.558
LLM-Prop-35M 1.353 0.074 2.160 2.007 0.104 685.379 400.681 1.706
Table 24: Results for Cantor HEA. The $\downarrow$ on the right of each property denotes the MAE error.
Input Model FEPA $\downarrow$ EPA $\downarrow$ Ehull $\downarrow$ VPA $\downarrow$ Wtd. Avg. (MAD/MAE) $\uparrow$
- Dummy (MAD) 0.108 0.792 0.175 0.470 -
Structure CGCNN (baseline) 0.012 0.016 0.018 0.164 17.780
Composition Llama 2-7b-chat:0S 21.986 8.097 61.275 15.079 0.034
Llama 2-7b-chat:5S 0.121 1.205 0.188 0.477 0.867
MatBERT-109M 0.033 0.049 0.034 0.189 6.769
LLM-Prop-35M 0.033 0.035 0.034 0.185 8.400
Structure Llama 2-7b-chat:0S 141.784 9.483 2791.970 107.467 0.022
Llama 2-7b-chat:5S Inval. Inval. Inval. Inval. Inval.
MatBERT-109M 0.015 0.045 0.019 0.081 9.958
LLM-Prop-35M 0.013 0.022 0.015 0.068 15.728
Description Llama 2-7b-chat:0S 212.607 7.843 1.062 44.740 0.069
Llama 2-7b-chat:5S 0.136 1.289 0.186 0.506 0.820
MatBERT-109M 0.015 0.045 0.019 0.080 9.976
LLM-Prop-35M 0.013 0.022 0.015 0.061 15.926
Table 25: Results for QMOF. The $\downarrow$ on the right of each property denotes the MAE error.
Input Model Bandgap $\downarrow$ Total energy $\downarrow$ LCD $\downarrow$ PLD $\downarrow$ Wtd. Avg. (MAD/MAE) $\uparrow$
- Dummy (MAD) 0.853 389.53 3.926 3.432 -
Structure CGCNN (baseline) 0.351 261.601 0.965 0.795 3.076
Composition Llama 2-7b-chat:0S 0.947 1498.974 87.894 395.609 0.303
Llama 2-7b-chat:5S 1.316 516.724 3.163 3.161 0.932
MatBERT-109M 0.468 229.838 1.686 1.461 2.049
LLM-Prop-35M 0.485 240.310 1.712 1.591 1.958
Structure Llama 2-7b-chat:0S 4.249 1598.672 191.713 323.684 0.119
Llama 2-7b-chat:5S Inval. Inval. Inval. Inval. Inval.
MatBERT-109M 0.428 88.984 1.350 1.218 3.024
LLM-Prop-35M 0.394 90.098 1.332 1.196 3.076
Description Llama 2-7b-chat:0S 2.384 1791.591 154.422 539.926 0.152
Llama 2-7b-chat:5S 1.098 546.695 3.491 2.933 0.946
MatBERT-109M 0.394 94.257 1.317 1.167 3.055
LLM-Prop-35M 0.408 90.342 1.387 1.213 3.016
Table 26: Results for JARVIS-QETB. The $\downarrow$ on the right of each property denotes the MAE error.
Input Model FEPA $\downarrow$ EPA $\downarrow$ Total energy $\downarrow$ Indirect Bandgap $\downarrow$ Wtd. Avg. (MAD/MAE) $\uparrow$
- Dummy (MAD) 1.261 719.975 88.432 0.210 -
Structure CGCNN (baseline) 0.642 3.155 7.883 0.038 61.729
Composition Llama 2-7b-chat:0S 438.919 1949.724 514.448 1.000 0.188
Llama 2-7b-chat:5S 1.553 694.574 85.674 0.161 1.047
MatBERT-109M 0.881 18.957 10.798 1.000 11.952
LLM-Prop-35M 0.443 3.400 4.150 0.113 59.443
Structure Llama 2-7b-chat:0S 489.087 1746.612 134.783 5.319 0.278
Llama 2-7b-chat:5S 1.576 703.03 82.224 0.123 1.152
MatBERT-109M 0.051 5.327 3.389 0.044 47.687
LLM-Prop-35M 0.054 2.262 1.835 0.114 97.919
Description Llama 2-7b-chat:0S 446.254 1763.805 182.817 1.316 0.264
Llama 2-7b-chat:5S 1.484 709.001 85.454 0.206 0.980
MatBERT-109M 0.048 5.859 3.007 0.027 46.586
LLM-Prop-35M 0.056 2.306 2.496 0.114 93.001
Table 27: Results for OQMD. The $\downarrow$ on the right of each property denotes the MAE error.
Input Model FEPA $\downarrow$ Bandgap $\downarrow$ Wtd. Avg. (MAD/MAE) $\uparrow$
- Dummy (MAD) 0.736 0.415 -
Structure CGCNN (baseline) 0.033 0.062 14.496
Composition Llama 2-7b-chat:0S 38.194 2.166 0.105
Llama 2-7b-chat:5S 0.726 0.318 1.160
MatBERT-109M 0.096 0.107 5.772
LLM-Prop-35M 0.080 0.146 6.020
Structure Llama 2-7b-chat:0S 80.815 8.767 0.028
Llama 2-7b-chat:5S 0.700 0.240 1.391
MatBERT-109M 0.053 0.058 10.521
LLM-Prop-35M 0.039 0.129 11.041
Description Llama 2-7b-chat:0S 29.677 2.221 0.106
Llama 2-7b-chat:5S 0.742 0.283 1.230
MatBERT-109M 0.049 0.059 11.027
LLM-Prop-35M 0.045 0.114 9.995
Table 28: Results for OMDB. The $\downarrow$ on the right of each property denotes the MAE error.
Input Model Bandgap $\downarrow$ MAD/MAE $\uparrow$
- Dummy (MAD) 0.803 -
Structure CGCNN (baseline) 0.292 2.751
Composition Llama 2-7b-chat:0S 0.907 0.885
Llama 2-7b-chat:5S 0.796 1.009
MatBERT-109M 0.517 1.554
LLM-Prop-35M 0.533 1.507
Structure Llama 2-7b-chat:0S 5.051 0.159
Llama 2-7b-chat:5S 0.864 0.930
MatBERT-109M 0.452 1.777
LLM-Prop-35M 0.452 1.777
Description Llama 2-7b-chat:0S 5.194 0.155
Llama 2-7b-chat:5S 0.802 1.001
MatBERT-109M 0.435 1.847
LLM-Prop-35M 0.485 1.656